### v2

@@ Abstract
Digital representations of individuals—“digital twins”—promise to transform social science and decision-making. Yet it remains unclear whether such twins truly mirror the people they emulate. We conducted 19 preregistered studies with a representative U.S. panel and their digital twins, each constructed from rich individual-level data, enabling direct comparisons between human and twin behavior across a wide range of domains and stimuli (including never-seen-before ones). Twins reproduced individual responses with 75 % accuracy and seemingly low correlation with human answers ($\approx$ 0.2). However, this apparently high accuracy was no higher than that achieved by generic personas based on demographics only. In contrast, correlation improved when twins incorporated detailed personal information, even outperforming traditional machine-learning benchmarks that require additional data. Twins exhibited systematic strengths and weaknesses—performing better in social and personality domains, but worse in political ones—and were more accurate for participants with higher education, higher income, and moderate political views and religious attendance. Together, these findings delineate both the promise and the current limits of digital twins: they capture some relative differences among individuals but not yet the unique judgments of specific people. All data and code are publicly available to support the further development and evaluation of digital twin pipelines.
keywords:
digital twins, large language models, social science experiments
1 Introduction
@@
Accurately measuring and predicting people’s opinions, preferences and reactions to interventions is fundamental to public policy, social sciences, and innovation. Yet, this essential practice has long been constrained by psychological, ethical, and practical barriers inherent to human subjects research. For example, psychological factors hinder repeated measures and counterfactual analyses; ethical concerns restrict high-stakes scenarios, harmful interventions, and access to rare populations; and practical barriers include the cost and logistical burden of recruiting participants.
@@
Large language models (LLMs) offer the potential to overcome these long-standing barriers through the scalable simulation of human responses [2, 7]. If such models are able to simulate the behavior of humans, they would not only offer faster and cheaper data collection, but potentially even “better” data for researchers and managers: LLMs do not tire, can be repeatedly surveyed, are capable of counterfactual responses (e.g., within-subject experiments in which the response to each stimulus is not influenced by previous stimuli), and can simulate high-stakes or potentially harmful scenarios without ethical risk. As a result, interest in simulating human responses using synthetic data has surged among academics from a wide range of disciplines spanning machine learning and social sciences, as well practitioners from a wide range of institutions spanning start-ups, polling companies, consulting firms and Fortune 500 companies.
Yet, there is currently a heated debate in the literature regarding the viability of leveraging LLMs to simulate human research. On the one hand, some have argued that LLMs are able to show human-like behavior (e.g., [12]) and to replicate many experimental treatment effects, at least directionally (e.g, [11, 6]). On the other, some have expressed skepticism, both on theoretical and empirical ground (e.g., [19, 8, 4, 15]). Studies have shown that LLMs often struggle to reproduce human opinions (e.g.,[18, 16, 14]) and that they tend to overestimate experimental effect sizes [6].
While most extant research in the area has relied on simulating the behavior of generic, synthetic personas defined primarily by demographic characteristics, one particularly promising approach relies on creating digital twins. Digital twins are created by combining (often via in-context learning, although RAG and fine-tuning are also possible) a base LLM (e.g., GPT, Llama) with rich individual-level information on a particular individual, in order to get the LLM to mimic that individual’s behavior. A panel of faithful digital twins could help improve the accuracy of synthetic data while better capturing heterogeneity across individuals. Beyond methodological advances, digital twins also open unique applications that distinguish them from other LLM simulations—from agentic AI (e.g., negotiating, working, or dating on one’s behalf) to self-reflection, where individuals interact with digital copies of themselves. Digital twins also complement recent work (e.g., [3, 13]) on fine-tuned LLMs trained on large cross-sectional datasets: these fine-tuned models may serve as alternative base LLMs for the creation of digital twins, i.e., they can be combined with additional panel data where participants are tracked across multiple experiments or studies, allowing the creation of twins associated with specific individuals. [^1] We have explored using Centaur [3] as a base model. Performance was overall not satisfactory. The relatively low performance may be due to limitations of the Llama base model on which Centaur is built or our formatting and out-of-distribution testing, among others. We leave a thorough investigation of the training of digital twins on such enhanced LLMs for future work. See SI for detail.
Despite their potential, to the best of our knowledge only two studies have examined digital twins to date. [17] created digital twins of over 1,000 individuals based on proprietary interview transcripts and reported encouraging results (relative accuracy of 85% on the General Social Survey, based on the ratio of digital twin accuracy to test-retest accuracy). [20] developed Twin-2K-500, a public dataset permitting the creation of digital twins of over 2,000 individuals based on their answers to over 500 questions. While they reported relatively high accuracy of digital twins on holdout data (average accuracy of 72%, relative accuracy of 88% based on the ratio of digital twin accuracy to test-retest accuracy), they found that digital twins replicated only about half of the between- and within-subject effects they tested. Notably, both of these studies focused on replicating known experiments and surveys, which may have been part of the training data of the base model.
This initial prior work leaves many questions unanswered: how well do digital twins mimic human behavior across performance metrics and domains, including new, never-seen-before studies? How do digital twins compare to relevant benchmarks such as generic personas based on demographic characteristics? How does digital twin performance vary across participant characteristics? across implementation methods (e.g., base LLM, temperature, use of fine-tuning)?
Answering these questions in a transparent and replicable manner calls for the creation of large, publicly available datasets that cover a wide range of domains. To this end, we conducted a mega-study consisting of 19 pre-registered (https://researchbox.org/4145) sub-studies run on the Twin-2K-500 human sample and their digital twins. Each digital twin was created based on answers provided by its human counterpart to 500 questions (approximately 128K characters) covering demographics (14 questions), personality traits (279 questions from 19 personality tests measuring 26 constructs), cognitive abilities (85 questions covering 11 measures), economic preferences (34 questions covering 10 measures), as well as responses to heuristics and biases experiments (48 questions from 17 experiments) and a pricing study (40 questions), collected during a four-wave study (average completion time of 145 min). See Figure 1 for an overview and SI for detail.
@@
Our 19 sub-studies were designed by a diverse group of co-authors, reflecting how scholars might employ digital twins in practice. They cover a wide range of domains and topics including creativity, politics, privacy preferences, storytelling, fairness perceptions, interactions with technology platforms, luxury consumption, news consumption, and labor market preferences, among others. See SI for details. Unlike prior work that primarily replicated published studies [17, 20], our sub-studies combined replications with unpublished research and newly designed studies, ranging from behavioral economics paradigms and personality scale development to correlational and within-subject designs. Together, they provide an ecologically valid test of digital twins as they are available today.
@@
Each sub-study was run on a sub-sample of the original Twin-2K-500 participants (recruited from Prolific) and their digital twins, i.e., we match the answer of each digital twin to each question to the actual answer given by their human counterpart to the same question. This enables us to test, across a wide range of domains, the accuracy of digital twins at the individual level as well as the correlation between responses from the digital twins and their human counterparts. Because our digital twins are based on a representative sample of the U.S. population, we can explore not only how performance varies across types of domains and questions, but also which groups are well represented by digital twins and which are not. We make our data (https://huggingface.co/datasets/LLM-Digital-Twin/Twin-2K-500-Mega-Study) and code (https://github.com/TianyiPeng/Twin-2K-500-Mega-Study) publicly available. Thus, our data also benefits other researchers and can provide a standardized testbed to help test and improve the performance of digital twins. [^2] Our data may be combined with the original Twin-2K-500 dataset, with twins matched by a Twin ID. For privacy reasons we are unable to share the Prolific ID associated with each Twin ID; therefore, other researchers are not able to run new studies on the human sample like we are.
We find that the individual-level accuracy of digital twins may at first appear high (approximately 75%), and the correlation between the answers from digital twins and their human counterparts may at first appear low (approximately 0.2). However, the individual-level accuracy achieved by digital twins is not significantly different from what is achieved with simpler, generic personas based on demographics only. On the other hand, correlation is improved by creating digital twins that leverage detailed individual-level information. Moreover, the correlation achieved by digital twins is higher than what may be achieved by traditional machine learning benchmarks that require additional data: matching the correlation performance of digital twins with an XGBoost model would require collecting the outcome of interest from (i.e., running the actual study with a training sample of) approximately 180 participants. On population-level metrics, we find that digital twins do not improve our ability to estimate the average response in the population compared to generic personas. While they improve our ability to capture variance, responses from digital twins remain under-dispersed. We also uncover specific domains in which digital twins tend to perform better (e.g., social domains, personality measures) and others in which they tend to perform worse (e.g., political domain). We find that digital twins tend to be more accurate for participants with higher education levels, higher income, and moderate political views and religious attendance habits. Finally, our best performing twins are build using in-context learning with GPT4.1 as base LLM and temperature of 0.
@@
Figure 1: Mega-Study Overview. We run 19 pre-registered studies on digital twins from the Twin-2K-500 dataset and their human counterparts. The studies were proposed by a diverse group of scholars and cover a wide range of behaviors and domains. As a set, they represent how digital twins may be leveraged today by social scientists. We match the answer of each digital twin to each question with the answer from their human counterpart, allowing us to explore the performance of digital twins both at the individual and population levels.
2 Results
2.1 Digital Twin Performance Across Metrics
2.1.1 Performance metrics
@@
We pre-registered 164 outcomes (e.g., willingness to share a particular article on social media, likelihood to apply for a particular job) across our 19 sub-studies. A key advantage of our study is the ability to directly match the response of each digital twin to that of their human counterpart. We focus on performance measures that leverage this feature of the data. Specifically, for each outcome, we match each twin’s response to its corresponding human’s response and compute two metrics:
@@
  • Individual-level accuracy: Following previous research [13, 20], we measure individual-level accuracy as 1-(MAD/range), where MAD is the mean absolute deviation between a human’s and their twin’s response, and range is the natural range of the outcome. [^3] We exclude outcomes without a defined range, such as open-ended price estimates, resulting in 161 observations for this metric. We compute the average individual-level accuracy across participants for each outcome, where a higher value indicates greater accuracy (between 0 and 1, where 1 denotes perfect accuracy).
@@
  • Correlation: We calculate the correlation between human and twin responses across participants for each outcome. This reflects the degree to which twins capture individual-level heterogeneity in human responses. Following previous research [17], we compute an overall average correlation across outcomes by first applying a Fisher z-transformation to individual correlations, averaging them, and then converting back using an inverse transformation.
@@
In addition to these metrics based on matched human-twin data, we compute two distribution-level performance metrics that compare the aggregate distributions of responses from twins and humans for each outcome:
@@
  • Comparison of averages: We compare the average response from humans to that of their twins for each outcome using the absolute value of Glass’s $\Delta$. [^4] Glass’s $\Delta$ is an effect size measure calculated as: $\Delta=(\hat{X}_{twin}-\hat{X}_{human})/s_{human}$, where $\hat{X}_{twin}$ and $\hat{X}_{human}$ are the sample means of the twin and human groups respectively, and $s_{human}$ is the standard deviation of the human group. It is similar to Cohen’s d, but it relies on the standard deviation from the human sample only, as the assumption of equal variances between the two samples is not valid in our case. Higher values indicate larger differences between group means (expressed in standard deviations).
@@
  • Comparison of standard deviations: We compute the ratio of standard deviations between humans and twins for each outcome. Values below (above) 1 suggest lower (higher) variance in twins compared to humans.
2.1.2 How do digital twins compare to their human counterparts?
See SI for histograms of all four performance metrics across outcomes. We first turn to our accuracy metric, which captures how well each digital twin predicted their human counterpart’s exact response. Across all outcomes, the average individual-level accuracy is 0.748, which compares favorably to the human test-retest accuracy of 0.817 reported by [20] using different tasks on the same sample of human respondents as the current study. However, we note that a random benchmark (selecting answers from a uniform distribution on the relevant range) achieves an individual accuracy level of 0.629, suggesting that the absolute levels of individual-level accuracy should be interpreted with caution. As an illustration, if the true responses were uniformly distributed on [0,1], always predicting 0 would yield an individual-level accuracy of 0.5. This does not imply that variations in individual-level accuracy are not meaningful, but this does stress the importance of comparing accuracy across relevant benchmarks, which we do later.
We next turn to our correlation metric, which captures how well twins reflect human-level heterogeneity. The correlation between twin and human responses is positive in 157 of the 164 outcomes (95.7%), and it is significantly positive with p<<0.05 in 128 of these cases (78.0%). The overall average correlation is 0.197.555Note that [17] report a higher correlation in their digital twin study. However, they compute the correlation across questions for each participant, while we compute the correlation across participants for each outcome.
Turning to distribution-level metrics, average twin responses differ from human responses by 0.352 standard deviations on average; [^6] This means the MSE of the digital twin’s mean estimate is about, $(0.352\sigma)^{2}\approx 0.12\sigma^{2}$ , which approximately equal to the MSE of the sample mean from 10 humans, $0.1\sigma^{2}$. a paired t-test finds a statistically significant difference (p<<0.05) between humans and digital twins in 125 out of 164 outcomes (76.2%). Finally, the standard deviation of the twin responses is lower than that of human responses in 154 of 164 cases (93.9%), indicating under-dispersion in twin responses (consistent with previous findings that responses simulated by LLMs tend to be less diverse than those generated by humans, e.g., [21]). This difference is statistically significant (p<<0.05) in 146 of those 154 cases.
2.1.3 How do digital twins compare to benchmarks that require less information?
The performance of digital twins is influenced both by the base LLM (GPT4.1 in this case) as well as the individual-level data used to create each twin. To assess the value of leveraging this individual-level data to create digital twins, we hold the base model constant and test how different levels of personal detail influence performance across metrics. The five benchmarks are (see SI for more detail):
  • Full Persona: complete data from the Twin-2K-500 dataset (approximately 128K characters), covering demographic, psychological, economic, personality and cognitive measures, as well as responses to classic behavioral economics experiments and a pricing survey. This is our default approach for creating digital twins, used throughout the paper unless otherwise specified.
  • Persona summary: a concise, statement-based summary (approximately 13K characters) of the questions and responses with distributional information (e.g., Big 5 personality scores with percentile ranks rather than 44 detailed questions and answers on which the scores are based).
  • Demographics only: includes the 14 demographic variables (region, sex, age, education, race, citizenship, marital status, religion, religious attendance, political party, household income, political ideology, household size, employment status) which are part of the Twin-2K-500 dataset (and therefore also part of the full persona and the persona summary). This benchmark captures the performance that would be obtained from generic personas based on demographic characteristics only.
  • Empty Persona: uses an identical system prompt for all twins (no individual information).
  • Random Responses: responses to each outcome are drawn from a uniform random distribution with a support equal to the outcome’s range (i.e., no LLM is required for this benchmark). [^7] For the three outcomes that do not have a natural range, we use the empirical range in humans’ answers as support for the random distribution.
Comparing the full persona version to the other versions shows that adding more detailed personal information improves correlation but has only a modest effect on individual-level accuracy (see Figure 2). [^8] We perform paired t-tests to compare each pair of benchmarks on each metric. For correlation we apply a z-transformation before conducting the t-test. We drop one outcome in these comparisons, the creativity rating of ideas generated by humans and their twins in one of the sub-studies, as it was provided by humans and was only available for ideas from twins based on the full persona approach. We also set correlation to 0 for cases where there is no variation among twin answers, which may happen with the “empty persona” or “demographics only” benchmarks. For example, twins built from full personas achieve only slightly higher individual-level accuracy than those based only on demographic information (p=0.37). While richer information increases the variance in twin responses, it does not increase the accuracy of their population mean estimate.
In other words, enriching digital twins with detailed individual-level information improves our ability to capture heterogeneity across participants and predict relative differences between them, that is, it enhances our ability to distinguish one participant from another. However, it does not substantially improve our ability to reproduce the exact responses given by specific participants, nor does it enhance predictions of population-level averages. We also see that the performance of twins is not statistically significantly different when the information is provided in raw vs. summary form for any of the four metrics.
@@ 2.1.3 How do digital twins compare to benchmarks that require less information?
Figure 2: Gains from Leveraging Individual-Level Data. *: best performing benchmark, or not significantly different from best at p<<0.05 (not applicable to ratio of standard deviations). Creating digital twins using rich individual-level data improves the correlation between twin and human responses compared to synthetic personas based on demographics only. It also increases the variance in responses, although digital twins remain under-dispersed. Individual-level accuracy and predictions of population means are not improved.

### v4

@@ Abstract
Scientists and practitioners are aggressively moving to deploy digital twins - LLM-based models of real individuals – across social science and policy research. We conducted 19 pre-registered studies with 164 diverse outcomes (e.g., attitudes towards hiring algorithms, intention to share misinformation) and compared human responses to those of their digital twins (trained on each person’s previous answers to over 500 questions). We find the digital twins’ answers are only modestly more accurate than those from the (homogeneous) base LLM and correlate weakly with human responses (average r = 0.20). We document five ways in which digital twins distort human behavior: (i) stereotyping, (ii) insufficient individuation, (iii) representation bias, (iv) ideological biases, (v) hyper-rationality. Together, our results caution against the premature deployment of twins, which may systematically misrepresent human cognition and undermine both scientific understanding and practical applications.
@@
Accurately measuring and predicting people’s opinions, preferences and reactions to interventions is fundamental to evidence-based policy, credible social science, and successful innovation. Yet, this essential practice has long been constrained by psychological, ethical, and practical barriers inherent to human subjects research. For example, contamination from prior exposure hinders repeated measures and counterfactual designs; the imperative to minimize harm restricts high-stakes scenarios, harmful interventions, and access to rare populations; and resource constraints, such as the cost and logistical burden of recruiting participants, limit the scope of measurement.
@@
Large language models (LLMs) offer the potential to overcome these long-standing barriers through the scalable simulation of human responses (?, ?, ?, ?). If such models are able to simulate human behavior, they would offer not only faster and cheaper data collection, but also potentially “better” data for researchers and managers: LLMs do not tire, can be repeatedly surveyed, and can simulate high-stakes or potentially harmful scenarios without ethical risk. As a result, interest in simulating human responses using synthetic data has surged among academics from a wide range of disciplines spanning machine learning and the social sciences, as well as practitioners from a wide range of institutions including start-ups (some of which are already valued at around \$1B [^1] TechCrunch, 12/05/2025.), polling companies, consulting firms and Fortune 500 companies.
Among the various tools producing synthetic data, digital twins are particularly promising. Digital twins are created by combining a base LLM (e.g., GPT, Llama) with rich individual-level information on a particular person, in order to get the LLM to mimic their specific behavior. If digital twins can faithfully represent individuals, they may address heterogeneity (?) and accuracy (?, ?, ?, ?) issues documented in prior synthetic data research. Beyond methodological advances, digital twins also open unique applications that distinguish them from other LLM simulations—from agentic AI (e.g., negotiating, working, or dating on one’s behalf) to self-reflection, where individuals interact with digital copies of themselves.
Two nascent studies of digital twins have shown promise. One created digital twins of over 1,000 individuals based on proprietary interview transcripts (?) and reported encouraging results (relative accuracy of 85% on the General Social Survey, based on the ratio of digital twin accuracy to test-retest accuracy). The other developed Twin-2K-500, a public dataset permitting the creation of digital twins of over 2,000 individuals based on their answers to over 500 questions (?), and reported relatively high accuracy of digital twins on holdout data (average accuracy of 72%, relative accuracy of 88% based on the ratio of digital twin accuracy to test-retest accuracy). This early promise coincides with increasing interest and deployment: the Twin-2K-500 dataset has been downloaded thousands of times, and more generally, digital twin applications are being frantically built, deployed and sold across academia and industry.
Yet, this rapid adoption of digital twins has outpaced rigorous empirical validation of whether they can actually serve as human surrogates. The existing evidence covered above, while encouraging, comes from a narrow subset of known experiments and surveys, which may overlap with the training data of the base model. This raises a fundamental question: can digital twins faithfully replicate the complexity of human cognition across domains and contexts, and, if not, does their widespread adoption introduce systematic distortions that threaten the validity of behavioral science? Answering these questions requires a transparent, rigorous, and comprehensive test using large, publicly available datasets that cover novel domains and tasks.
To that end, we quantify digital twin performance using a mega-study consisting of 19 pre-registered (https://researchbox.org/4145) sub-studies covering 164 diverse outcomes, conducted on the Twin-2K-500 human sample and their corresponding digital twins (cumulative sample size of N=13,506 human participants across sub-studies, 1,784 unique participants). Our evaluation demonstrates underwhelming performance overall, which we unpack by documenting five ways in which digital twins distort human behavior: (i) stereotyping, (ii) insufficient individuation, (iii) representation bias, (iv) ideological biases, (v) hyper-rationality. Our results suggest that the premature deployment of twins risks systematically misrepresenting human cognition in ways that could undermine both scientific understanding and practical applications, while also pointing to future avenues for improvement.
Empirical Context
@@
Our diverse group of co-authors designed 19 sub-studies to reflect how scholars might employ digital twins in practice. These studies cover a wide range of domains and topics including creativity, politics, privacy preferences, storytelling, fairness perceptions, interactions with technology platforms, luxury consumption, news consumption, and labor market preferences, among others. Unlike prior work that primarily examined digital twins’ performance in published studies (?, ?), our sub-studies combined established paradigms from published research, new paradigms from unpublished research, and newly designed studies. Our contexts range from behavioral economics paradigms to personality scale development and include correlational, within-subject, and between-subject designs. Together, they span a considerably broader and richer scope of use cases than any paper to date, providing an ecologically valid test of digital twins as they are available today.
To create each digital twin, we used in-context learning based on answers to over 500 questions (approximately 128K characters) provided by the twin’s human counterpart. These questions covered demographics (14 questions), personality traits (279 questions from 19 personality tests measuring 26 constructs), cognitive abilities (85 questions covering 11 measures), economic preferences (34 questions covering 10 measures), as well as heuristics and biases experiments (48 questions from 16 experiments), and a pricing study (40 questions), collected during a four-wave, longitudinal study (average completion time of 145 min). These (human) data have high test-retest accuracy, good face validity, and largely replicate known effects, suggesting they provide high-quality training data for digital twins (?).
@@
In each sub-study, both human participants and their corresponding digital twins answered the exact same questions. By matching each digital’s answers to those given by its human counterpart to the same questions, we are able to test, across a wide range of domains, the individual-level accuracy of digital twins as well as the correlation between responses from the digital twins and their human counterparts. Further, because our digital twins are based on a representative sample of the U.S. population, we can explore not only how performance varies across types of domains and questions, but also which groups are well represented by digital twins and which are not. We share the detailed analyses of each sub-study for researchers interested in using digital twins in specific domains (see SI). We make our data (https://huggingface.co/datasets/LLM-Digital-Twin/Twin-2K-500-Mega-Study) and code (https://github.com/TianyiPeng/Twin-2K-500-Mega-Study) publicly available. [^2] Researchers may combine our data with the original Twin-2K-500 dataset, with twins matched using a Twin ID. For privacy reasons we are unable to share the Prolific ID associated with each Twin ID; therefore, other researchers are not able to run new studies on the human sample as we did. See Figure 1 for an overview and SI for details of our mega-study.
@@
Figure 1: Mega-Study Overview. We ran 19 pre-registered studies on digital twins from the Twin-2K-500 dataset and their human counterparts. The studies were proposed by a diverse group of scholars and cover a wide range of behaviors and domains. As a set, they represent how digital twins may be leveraged today by social scientists. We match the answer of each digital twin to each question with the answer from their human counterpart, allowing us to explore the performance of digital twins at both the individual and population levels.
Results
Digital Twin Performance
Performance metrics
@@ Results
We pre-registered 164 outcomes (e.g., willingness to share a particular article on social media, likelihood of applying for a particular job) across our 19 sub-studies. We focus on performance measures that directly match the response of each digital twin to that of its human counterpart. Specifically, for each outcome, we match each twin’s response to that of its corresponding human, and compute two metrics:
@@
  • Individual-level accuracy: Following previous research (?, ?), we measure individual-level accuracy as 1-(MAD/range), where MAD is the mean absolute deviation between a human’s response and that of their twin, and range is the natural range of the outcome. [^3] We exclude outcomes without a defined range, such as open-ended price estimates, resulting in 161 outcomes for this metric. We compute the average individual-level accuracy across participants for each outcome, where higher values indicate greater accuracy (between 0 and 1, where 1 denotes perfect accuracy). For example, on a 1 to 7 scale, if the human answered 5 and the twin answered 6, accuracy would be 0.833.
@@
  • Correlation: We calculate the correlation between human and twin responses across participants, for each outcome. This metric reflects the degree to which twins capture individual-level heterogeneity in human responses. Following previous research (?), we compute an overall average correlation across outcomes by first applying a Fisher z-transformation to individual correlations, averaging these values, and then converting back using an inverse transformation.
@@
In addition to the metrics above, which are based on matched human-twin data, we also compute two distribution-level performance metrics that compare the aggregate distributions of responses from twins and humans for each outcome. These metrics are more commonly available (no individual-level matching required), but arguably less informative:
@@
  • Comparison of averages: We compare the average response of humans to that of their twins for each outcome using the absolute value of Glass’s $\Delta$. [^4] Glass’s $\Delta$ is an effect size measure calculated as: $\Delta=(\hat{X}_{twin}-\hat{X}_{human})/s_{human}$, where $\hat{X}_{twin}$ and $\hat{X}_{human}$ are the sample means of the twin and human groups respectively, and $s_{human}$ is the standard deviation of the human group. It is similar to Cohen’s d, but it relies on the standard deviation from the human sample only, as the assumption of equal variances between the two samples is not valid in our setting. Higher values of this metric indicate larger differences between group means (expressed in standard deviations).
@@
  • Comparison of standard deviations: We compute the ratio of standard deviations between humans and twins for each outcome. Values below (above) 1 suggest lower (higher) variance in twins relative to humans.
Performance comparisons
By default (and unless otherwise specified), we build each twin using a “full persona” approach which augments the LLM’s context window with that human’s complete data from the Twin-2K-500 dataset. Across all outcomes, the average individual-level accuracy of the “full persona” approach is 0.748. See Figure 2, and the SI for histograms of this and other performance metrics across outcomes for the “full persona” approach.
@@
Figure 2: Gains from Leveraging Individual-Level Data. *: best performing benchmark, or not significantly different from best at p<<0.05 (not applicable to ratio of standard deviations). Creating digital twins using rich individual-level data leads to a modest improvement in accuracy over predictions made without any individual-level data (empty persona) and no significant improvement over personas based on demographics only. Full personas improve the correlation between twin and human responses compared to these benchmarks, although the correlation is modest and digital twins remain under-dispersed.
To contextualize this level of accuracy and quantify gains from leveraging individual-level data for out-of-distribution predictions, we compare this performance to several benchmarks. First, we compare performance to that of random responses, where responses to each outcome are drawn from a uniform random distribution with support equal to the outcome’s range (i.e., no LLM is required for this benchmark). The random benchmark already achieves an individual accuracy level of 0.629, suggesting that absolute levels of individual-level accuracy should be interpreted with caution. As an illustration, if the true responses were uniformly distributed on [0,1], always predicting 0 would yield an individual-level accuracy of 0.5.
Next, we compare the “full persona” approach with a simple “empty persona” approach, which uses an identical prompt for all respondents (i.e., no individual information) and thus only reflects the base model’s behavior (with serving status and characteristics of the base model such as temperature being the only source of variation). The empty persona benchmark achieves an average accuracy of 0.734, which is statistically significantly, but only modestly, lower than full personas (a difference of 0.014). [^5] We perform paired t-tests to compare each pair of benchmarks on each metric. We drop one outcome in these comparisons, the creativity rating of ideas generated by humans and their twins in one of the sub-studies, as it was provided by humans and was only available for ideas from twins based on the full persona approach.
We also introduce a “demographics only” benchmark, which uses a persona description that only includes the 14 demographic variables (region, sex, age, education, race, citizenship, marital status, religion, religious attendance, political party, household income, political ideology, household size, employment status) which are part of the Twin-2K-500 dataset (and therefore also included in the full persona description). This benchmark captures the performance that would be obtained from generic personas based on demographic characteristics only. We find that full-persona twins achieve only slightly higher individual-level accuracy than those based only on demographic information ($Accuracy_{full\ persona}=0.748$, $Accuracy_{demographics\ only}=0.746$, $p=0.37$). This provides initial evidence that, when predicting an individual’s response, digital twins may mostly rely on stereotypical demographic-based tendencies rather than modeling an individual’s cognition. We explore this phenomenon and the comparison between full-persona and demographics-only twins more thoroughly in a subsequent section.
At the aggregate level, average twin responses differ from human responses by 0.352 standard deviations on average; [^6] This means the MSE of the digital twin’s mean estimate is about $(0.352\sigma)^{2}\approx 0.12\sigma^{2}$, which is approximately equal to the MSE of the sample mean from 10 humans, $0.1\sigma^{2}$. a paired t-test indicates a statistically significant difference (p<<0.05) between humans and digital twins in 125 out of 164 outcomes (76.2%).
We next turn to our correlation metric, which captures how well twins reflect human-level differences in individual responses across participants. The correlation between twin and human responses is positive in 157 of the 164 outcomes (95.7%), and significantly positive ($p<0.05$) in 128 of these cases (78.0%). The overall average correlation is 0.197.777Note that (?) report a higher correlation in their digital twin study. However, they compute the correlation across questions for each participant, while we compute the correlation across participants for each outcome, which is more often the measure of interest. On that metric, full personas perform markedly better than demographics-only personas (r=0.145), empty personas (r=0.080) and random responses (r=0.001). [^8] For correlation, we apply a z-transformation before conducting t-tests. We set correlation to 0 for cases where there is no variation among twin answers, which may occur with the “empty persona” or “demographics only” benchmarks. For the three outcomes that do not have a natural range, we use the empirical range in humans’ answers as support for the random distribution. The correlation achieved by empty personas is not 0 due to variations in stimuli within an outcome across participants, e.g., different respondents may see different versions of the same question or may be exposed to various conditions. This creates meaningful variations across respondents, even if predictions are not customized based on individual characteristics. Nevertheless, the correlation achieved by digital twins is modest by most social science standards, and comparable to the correlation found between height and intelligence (?).
Put together, enriching digital twins with detailed individual-level information significantly improves individual-level prediction, that is, the ability to reproduce the exact responses given by specific participants. However, the improvement over an empty persona is arguably negligible. Second, digital twins significantly improve our ability to capture heterogeneity across participants and predict relative differences between them, that is, it enhances our ability to distinguish one participant from another. However, the correlation between answers from digital twins and their human counterparts is modest.
Digital Twin Performance Across Implementation Methods
All results reported so far use GPT4.1 (dated 2025-04-14) with a default temperature of 0.7. One might wonder whether poor digital twin performance was in part due the specific configurations we employed rather than fundamental limitations of current digital twin technology. To explore the role of these factors, we systematically varied the temperature and base LLM (GPT-5, Deepseek, Gemini - including the latest Gemini-3 Pro model - and a version of GPT4.1 fine-tuned on the Twin-2K-500 dataset). We find the best results overall using GPT4.1 with temperature = 0 (e.g., highest correlation of 0.232 across all implementation options, tied for best on individual-level accuracy at 0.752). We also experimented with replacing the full persona information with a concise (approximately 13K characters), statement-based summary of the questions and responses with distributional information (e.g., Big 5 personality scores with percentile ranks rather than 44 detailed questions and answers on which the scores are based). We find that this simpler version performs very similarly to the full persona, offering a viable lower-cost alternative. See SI for details.
We also explored using an LLM fine-tuned specifically for human behavioral prediction as the base model for the creation of digital twins (?, ?). These fine-tuned models, trained on large cross-sectional datasets, may indeed serve as alternative base LLMs, i.e., they can be combined with additional panel data in which participants are tracked across multiple experiments or studies, allowing the creation of twins associated with specific individuals. We explored using Centaur (?) as a base model. Performance was lower than that achieved using GPT as the base model. The relatively low performance may be due to limitations of the Llama base model on which Centaur is built, or to our formatting and out-of-distribution testing, among other reasons. We leave a thorough investigation of the training of digital twins on such enhanced LLMs for future work. See SI for details.
Overall, twin performance across temperatures, base models, and persona formatting was, at best, modest.
Five Key Distortions
We next characterize systematic distortions in digital twins’ representations of human behavior that may underlie their poor performance. This is an important step toward better understanding how digital twins differ from humans, what implications their premature deployment might have, and how digital twin pipelines may be improved in the future. In particular, we highlight five key distortions.
1. Stereotyping
One possibility for digital twins’ modest performance is that their predictions do not sufficiently integrate the rich information on each individual’s psychological and cognitive profiles, and instead over-rely on generic characteristics such as demographics. The fact that full-persona twins achieve only slightly higher individual-level accuracy than those based only on demographic information provides initial evidence for stereotyping.
To test this possibility more directly, we compute the mean absolute deviation (MAD) between the answers (normalized between 0 and 1) given by full-persona twins vs. demographics-only twins. The average MAD across outcomes is 0.132, which is significantly lower than the average MAD between full-persona twins and empty-persona twins ($MAD_{full\ vs.\ demographics}=0.132$, $MAD_{full\ vs.\ empty}=0.175$, $p<0.01$) or the actual human answers ($MAD_{full\ vs.\ demographics}=0.132$, $MAD_{full\ vs.\ humans}=0.252$, $p<0.01$). That is, answers from full-persona twins are closer to those from demographics-only twins than to those from empty-persona twins or real humans.
As discussed earlier, correlation on the other hand improves more substantially when using full vs. demographics-only personas. This suggests that, to the extent full-persona twins deviate from demographics-only twins, they do so in a way that is somewhat consistent with variations across humans.